# Reassessing Noise Augmentation Methods

### v3

@@ 6 Discussion and Conclusion
Our findings reveal that across the board, ASR models trained with noise augmentation – encompassing background noise, speed variations – and reverberations exhibit enhanced robustness against adversarial attacks compared to those without such augmentation. Interestingly, the implementation of speed variations alone as a form of augmentation also contributed positively to the robustness of the models, albeit to a lesser extent than the combination of multiple noise types. This hints that robustness gains may be attributed to the increased diversity and volume of the dataset. Through our investigation, it becomes evident that incorporating speed and noise augmentation into the training pipeline is beneficial for both noise robustness and adversarial robustness.
@@
The findings from our study suggest several intriguing directions for further research. First, while noise augmentation has proven effective, understanding the optimal combination and intensity of different augmentation types remains an open question. The relative contributions of background noise, speed variations, and reverberations need to be quantified to develop more targeted augmentation strategies. Additionally, the role of dataset size and diversity in enhancing robustness merits deeper exploration. Our results indicate that simply increasing the dataset size with varied augmentations can lead to significant robustness improvements. Future studies could systematically vary the size and diversity of training data to better understand this relationship. Moreover, our research highlights the complex interplay between different forms of augmentation and their impact on adversarial robustness. A thorough investigation into how these augmentations interact with various attack types could provide more granular insights into defense mechanisms.
@@
For future work, we believe that exploring the interplay between noise augmentation and other defense strategies, such as adversarial training and model ensembling, could yield a more robust and multi-faceted approach to securing ASR systems against a wide array of threats. Combining these strategies could leverage the strengths of each, potentially leading to ASR systems that are not only robust against noise and adversarial attacks but also adaptable to evolving threats in real-world applications. By continuing to refine augmentation techniques and exploring their integration with other defensive measures, we can move towards developing ASR systems that are both resilient and reliable in the face of diverse and sophisticated adversarial challenges.

### v4

@@ 6 Discussion and Future Work
Our findings reveal that across the board, ASR models trained with noise augmentation – encompassing background noise, speed variations, and reverberations – exhibit enhanced robustness against adversarial attacks compared to those without such augmentation. Interestingly, the implementation of speed variations alone as a form of augmentation also contributed positively to the robustness of the models, albeit to a lesser extent than the combination of multiple noise types. This hints that robustness gains may be attributed to the increased diversity and volume of the dataset. Through our investigation, it becomes evident that incorporating speed and noise augmentation into the training pipeline is beneficial for both noise robustness and adversarial robustness.
@@
The findings from our study suggest several intriguing directions for further research. First, while noise augmentation has proven effective, understanding the optimal combination and intensity of different augmentation types remains an open question. The relative contributions of background noise, speed variations, and reverberations need to be quantified to develop more targeted augmentation strategies. Additionally, the role of dataset size and diversity in enhancing robustness merits deeper exploration. Our results indicate that simply increasing the dataset size with varied augmentations can lead to significant robustness improvements. Future studies could systematically vary the size and diversity of training data to better understand this relationship. In particular, this should encompass varying levels of signal-to-noise ratio and different noise dataset sizes. Moreover, our research highlights the complex interplay between different forms of augmentation and their impact on adversarial robustness and noise robustness. In particular, it could be relevant to study what impact noise-augmented training has on clean speech quality or which artifacts may be introduced by noise-augmented training (and potentially exploited by adversarial attacks). A thorough investigation into how these augmentations interact with the robustness w.r.t. various attack types could provide more granular insights into defense mechanisms.
@@
For future work, we believe that exploring the interplay between noise augmentation and other defense strategies, such as adversarial training and model ensembling, could yield a more robust and multi-faceted approach to securing ASR systems against adversarial attacks. By continuing to refine augmentation techniques and exploring their integration with other defensive measures, we can move towards developing ASR systems that are both robust and reliable in the face of sophisticated adversarial challenges.
7 Conclusion
In conclusion, our study provides a foundational understanding of the benefits of noise augmentation for ASR systems. Our comprehensive study has shed light on the efficacy of noise augmentation as a strategy for improving the robustness of automatic speech recognition systems with respect to noise as well as to adversarial examples. By conducting a systematic analysis of four ASR architectures, each trained with different augmentation techniques and subjected to a spectrum of adversarial attacks, we have provided empirical evidence supporting the benefits of noise augmentation for adversarial robustness.
@@ 8 Acknowledgements
Appendix A Noise Augmentation in Training
The following code snippet displays the corruptions used for the non-augmented model.
1env_corrupt: !new:speechbrain.lobes.augment.EnvCorrupt
2 openrir_folder: <folder>
3 babble_prob: 0.0
4 reverb_prob: 0.0
5 noise_prob: 0.0
6 noise_snr_low: 0
7 noise_snr_high: 15
8
9augmentation: !new:speechbrain.lobes.augment.TimeDomainSpecAugment
10 sample_rate: 16000
11 speeds: [100]
The following code snippet displays the corruptions used for the speed-augmented model.
1env_corrupt: !new:speechbrain.lobes.augment.EnvCorrupt
2 openrir_folder: <folder>
3 babble_prob: 0.0
4 reverb_prob: 0.0
5 noise_prob: 0.0
6 noise_snr_low: 0
7 noise_snr_high: 15
8
9augmentation: !new:speechbrain.lobes.augment.TimeDomainSpecAugment
10 sample_rate: 16000
11 speeds: [95, 100, 105]
The following code snippet displays the corruptions used for the fully augmented model.
1env_corrupt: !new:speechbrain.lobes.augment.EnvCorrupt
2 openrir_folder: <folder>
3 babble_prob: 0.1
4 reverb_prob: 0.1
5 noise_prob: 1.0
6 noise_snr_low: 0
7 noise_snr_high: 15
8
9augmentation: !new:speechbrain.lobes.augment.TimeDomainSpecAugment
10 sample_rate: 16000
11 speeds: [95, 100, 105]
Appendix B Noise Augmentation in Testing
The following code snippet displays the corruptions used in the augmentation for testing robustness.
1env_corrupt: !new:speechbrain.lobes.augment.EnvCorrupt
2 openrir_folder: <folder>
3 babble_prob: 0.0
4 reverb_prob: 0.0
5 noise_prob: 1.0
6 noise_snr_low: 0
7 noise_snr_high: 15
8augmentation: !new:speechbrain.lobes.augment.TimeDomainSpecAugment
9 sample_rate: 16000
10 speeds: [100]